OpenStudy (anonymous):

Solve the formula P = 2l + 2w for w.

OpenStudy (anonymous):

what do you mean to solve it?

OpenStudy (anonymous):

do u mean making w the subject of formula?

OpenStudy (anonymous):

2w = P-2l therefore w = (P-2l)/2

OpenStudy (anonymous):

which answer would it be a. w=P-21 B.w = p/21 c.w =2p/21 d. w= p-21/2

OpenStudy (anonymous):

Lumenaire is right your answer is part d dear!

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thank yaa